a number of espresso coffee stands in busy suburban malls. The foxed weekly expense of a o tand is $1,000 and the variable cost per oup of coffee served is $0.34. Required: 1. Fill in the following table with your estimates of total costs and cost served to 3 decimal places.) per cup of coffe at the indicated levels of activity for a coffee stand. (Round the Average cost per cup of cotf Cups of Coffee Served in a Week 1,400 1,500 Fixed cost Variable cost Total cost cost per cup of coffee of ooffee served increase, decrease, or remain the same as the number of cups of coffee served in a week increases? O Increase O Decrease Remain the same
Answer #1
1400 1500 1600
Fixed cost $1000 $1000 $1000
Variable cost (1400*0.34)=$476 (1500*0.34)=$510 (1600*0.34)=$544
Total cost (1000+476)=$1476 (1000+510)=$1510 (1000+544)=$1544
Average cost per cup of coffee served(Total cost/Cups of coffee served) (1476/1400)=$1.054 (1510/1500)=$1.007 (1544/1600)=$0.965

Hence average cost per cup of coffee served decreases as the number of cups of coffee served increases.

NOTE:Total fixed cost and variable cost per unit do not change with change in units.